OpenStudy (anonymous):

(x-y)^3 over y^3+z^3 divided by x^2-2xy+y^2 over y^2-yz+z^2

sam (.sam.):

\[x^{3}-xyz+xz ^{2}-x ^{2}y+y ^{2}z-yz ^{2}\]------------------------------- \[y ^{3}+z ^{3}\]

OpenStudy (anonymous):

\[\frac{(((x-y){}^{\wedge}3) / (y{}^{\wedge}3+z{}^{\wedge}3))}{(x{}^{\wedge}2-2x y+y{}^{\wedge}2)/(y{}^{\wedge}2-y z+z{}^{\wedge}2)}\text{=}\frac{x-y}{y+z} \]